How much work is done to bring two electrons to a distance of 5.3 x 10-11 m to the nucleus of a Helium atom (q = 3.2 x 10-19 C)?
1
Expert's answer
2020-10-21T09:52:06-0400

Solution

Work done is given by

"W=-\Delta U=\frac{kq_1q}{r}"

Where q1 is charge of electron.

q is charge of helium.

So work done is given

"W=\frac{9\times10^{9}\times2\times1.6\times10^{-19} \times3.2 \times10^{-19}}{5.3 \times10^{-11} }"

"W=17.39\times 10^{-18} J"